OpenStudy (love_to_love_you):

The sum of two numbers is 8. Find the smallest possible value for the sum of the cube of one number and the square of the other.

OpenStudy (mathmale):

Please begin by representing these two numbers. You could choose the letters x and y to do that. Using these letters, write an equation for "the sum of two numbers is 8." Choose one of the numbers (doesn't matter whether you choose x or y). Square it. Choose the other number. Cube it. Add together the square and the cube. Using your first equation, eliminate one of the two variables. Assuming that you know calculus: take the first deriv. of this sum function and set the deriv. = to 0. Solve the resulting eq'n for whichever variable is left (x or y). Know what to do next? Check your results.

OpenStudy (love_to_love_you):

x^2+(-x)^3=8

OpenStudy (love_to_love_you):

So I got to this point

OpenStudy (love_to_love_you):

x = 1.716189

OpenStudy (love_to_love_you):

Idk how to do the derivative stuff, I don't know very much calculus. It's all very blurry

OpenStudy (mathmale):

But what about that '8'? The sum of your two numbers is 8. You'll need to use this fact to eliminate one of the variables. Sum is 8: x+y=8. Solve this for either x or y, please.

OpenStudy (love_to_love_you):

y= 0x + 8

OpenStudy (love_to_love_you):

-x

OpenStudy (mathmale):

The 'sum of x and y' would be x+y. Since the sum of x and y is 8, we can write x+y=8. You cannot eliminate the '8', nor can you eliminate both variables. Again I give you x+y=8. Solve this for x (altho' you could solve for y as an alternative).

OpenStudy (love_to_love_you):

x = -y + 8

OpenStudy (mathmale):

Yes, that's right. You subtracted y from both sides. What is the square of x, if x=-y+8? Here you are "squaring a binomial."

OpenStudy (mathmale):

x^2=?

OpenStudy (mathmale):

Square of the sum of two numbers a and b is (a+b)^2=a^2+2ab+b^2. Square of the difference of two numbers a and b is (a-b)^2 = a^2 - 2ab -b^2.

OpenStudy (mathmale):

Thus, the square of (8-y) is what?

OpenStudy (love_to_love_you):

(8-y)^2 = 8^2 - 28y - y^2?

OpenStudy (mathmale):

Actually, the middle term is -16y. How did you get -28y?

OpenStudy (love_to_love_you):

2(8y)

OpenStudy (love_to_love_you):

I was just typing it out, I hadn't done any of the math yet

OpenStudy (mathmale):

2*8=16, right? I see. Can't write "28" for "2 times 8." Must write 2(8) or 2*8.

OpenStudy (love_to_love_you):

yes

OpenStudy (mathmale):

"The sum of two numbers is 8. Find the smallest possible value for the sum of the cube of one number and the square of the other." x+y=8, so x=-y+8. Correct. Thus, x^2=(-y+8)^2 = y^2 -16y +64. This is the "square of one number," and that number happens to be x. In terms of y, that number happens to be y^2 -16y +64. OK with that? "square of one number"

OpenStudy (mathmale):

We are dealing exclusively with 'y,' having solved x+y=8 for x. OK with that? Square of one number (in terms of y) is y^2-16y+64. Cube of one number (y) is y^3. Write out the sum of "the square of one number and the cube of the other."

OpenStudy (love_to_love_you):

(y^2 - 16y + 65) + (y^3)

OpenStudy (mathmale):

That's 64, not 65. Otherwise your expression is fine. That's "the sum of the square of one number and the cube of the other." At this point we have to decide what to do next. If you're able to find the derivative of this function of yours, do it. If you're not, then you could graph this function on a graphing calculator. You would see right away where the minimum value is and approximate the y value there. Your choice?

OpenStudy (love_to_love_you):

sorry, typo I'll graph it

OpenStudy (love_to_love_you):

or how could I find the derivative?

OpenStudy (mathmale):

Before I explain that I need to ask just how long you have studied calculus. Do you remember the "power rule?"

OpenStudy (love_to_love_you):

Is it when you take the exponent, multiply it by the coefficient and subtract 1 from the exponent?

OpenStudy (mathmale):

Yes. What is the derivative with respect to y of y^3? of y^2?

OpenStudy (love_to_love_you):

eeek, okay I understand it but I might get it wrong 3y^2?

OpenStudy (love_to_love_you):

2y?

OpenStudy (mathmale):

yes. yes. It's very important that you review the power rule in differentiation. if f(y)=y^3 + y^2 - 16y + 64, then the derivative will be f '(y) = 3y^2 - 2y - 16. Set that = to 0 now. Are you able to solve this quadratic equation for y?

OpenStudy (love_to_love_you):

y = 8/3, -2 ?

OpenStudy (mathmale):

Beg your pardon. That derivative is f '(y) = 3y^2 + 2y - 16. One root is y=+2. That agrees perfectly with my graph of f(y)= y^3 + y^2 - 16y + 64. This graph shows a minimum at y=2.

OpenStudy (mathmale):

Any questions about this process? x+y=8, x=-y+8 x^2 = y^2 -16y + 64 y^2 + y^3 = y^2 - 16y + 64 + y^2

OpenStudy (mathmale):

Finding the derivative of that last function results in 3y^2 +2y -16. We set this = to 0 and solve for y. y=2. OK?

OpenStudy (mathmale):

How would you find "the smallest possible value of the sum of the square of one number and the cube of the other?"

OpenStudy (mathmale):

Hint: You already have a function f(y) for this. It is y^3 + y^2 - 16y + 64.

OpenStudy (love_to_love_you):

16? by plugging in y?

OpenStudy (mathmale):

Yes, plug in y=2: let y=2 in y^3 + y^2 - 16y + 64.

OpenStudy (love_to_love_you):

44?

OpenStudy (mathmale):

Perfect. You know your stuff, really. That's your answer.

OpenStudy (mathmale):

y=2. You must also find x. Remember how x and y are related?

OpenStudy (love_to_love_you):

x = 6

OpenStudy (mathmale):

perfect. I've just given you a medal.

OpenStudy (love_to_love_you):

Thank you for the help! You're the best!

OpenStudy (mathmale):

Thank you so much. I love to love you. ;) Bye!

OpenStudy (love_to_love_you):

Haha bye :)